OpenStudy (anonymous):

A target consists of five concentric circles, each with a radius 1 inch longer than that of the previous circle. Calculate the area between the third and fourth circles from the center.

OpenStudy (anonymous):

does the first circle have radius 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the radius of the first circle is 1

OpenStudy (anonymous):

then the fourth circle has radius 4 and the third circle has radius 3 area is \[\pi r^2\] in general so fourth circle has area \[16\pi\] and third has area \[9\pi\] and the difference is \[16\pi-9\pi=7\pi\]

OpenStudy (anonymous):

YESSSS!!!!YOU GOT IT!!!....thanks so much...i have some more questions that i am going to post